# AdvDiff: Generating Unrestricted Adversarial Examples using Diffusion Models

### v3

@@ 5.2 Generation Quality: True ASR for UAEs
FID score is calculated by $d^{2}((m,C),(m_{w},C_{w}))=||m-m_{w}||^{2}_{2}+Tr(C+C_{w}-2(CC_{w})^{1/2})$, where $(m,C)$ measures the Gaussian distribution of the data in feature space. Because both U-GAN and AdvDiff do not modify the parameter of the generative model, we are expected to witness similar FID scores. However, IS score is $\text{exp}(\mathbb{E}_{x}\textbf{KL}(p(y|x)||p(y)))$. The conditional label distribution $p(y|x)$ should have low entropy as the adversarial objective aims to fool the classifier with high confidence on the target attack label, while the $p(y)$ should be a uniform distribution. As we receive both high ASR on the ImageNet dataset with U-GAN and AdvDIff, the reason for the significant drop in U-GAN is the change of distribution $p(y)$. In other words, the U-GAN’s UAEs are not equally distributed among 1000 classes, and the most possible reason is the frequent generation of flipped-label UAEs and low-quality UAEs, which can also be inferred by the drop in Precision. Figure 4 illustrates this problem with some examples.
We further observed an obvious increase in Recall by U-GAN with BigGAN, which indicates that U-GAN’s $z_{adv}$ fails to follow the normal distribution and often has a high variance, which covers a larger latent space and increases the Recall. However, most state-of-the-art GANs [2, 18] adopt truncation tricks to GAN latents to improve the generation quality, which reduces the diversity (Recall). Therefore, GAN-based UAEs perform even worse with better GAN networks. In conclusion, evaluating generation quality is critical to measure the attack performance of unrestricted adversarial attacks, especially the IS score. UAEs with much lower generation performance can be easily detected by humans.
5.3 UAEs against Defenses and Black-box Models
Experiments on MNIST and ImageNet datasets show that adversarial defense methods like adversarial denoising [26] and adversarial training [24] fail to defend against UAEs. Current defenses assume the adversarial examples are based on perturbations over data from the training dataset, i.e., $x_{adv}=x+\nabla\mathcal{L},x\in D$. However, UAEs are synthetic data generated by the generative model. Because of different data sources, current defenses are hard to defend UAEs, which brings severe security concerns to deep learning applications. We also test the attack transferability of AdvDiff and the results show that it achieves a moderate performance against black-box models. Experiment results are given in Table 5.
  Method ResNet-152 [13] Inception v3 [38] DenseNet-121 [17]
PGD 31.2 14.7 36.0
U-BigGAN 18.4 22.1 16.8
AdvDiff-PGD 26.3 24.9 38.6
AdvDiff 20.5 10.9 35.8
  Method MobileNet v2 [33] PNASNet [23] MNASNet [39]
PGD 38.4 11.4 35.6
U-BigGAN 14.2 8.5 16.1
AdvDiff-PGD 22.4 22.1 37.1
AdvDiff 15.4 23.2 38.9
  Method Adv-Inception [24] AdvProp [42] DiffPure [26]
PGD 38.4 11.4 35.6
U-BigGAN 14.2 8.5 16.1
AdvDiff-PGD 25.1 89.0 98.0
AdvDiff 16.1 89.3 98.1
5.4 Better Adversarial Diffusion Sampling
We present a detailed comparison with PGD-based adversarial guidance [5, 4, 6] in Table 3 and 4. The results show that the proposed adversarial guidance achieves significantly higher generation quality than PGD-based adversarial guidance. With PGD gradient guidance, the diffusion model generates images with a similar Recall score but a much lower Precision score, which indicates that the PGD gradient influences the benign reverse generation process and causes the generation of low-quality images. The result proves that the adversarial guidance of diffusion models should be carefully designed without affecting the benign sampling process. However, PGD-based AdvDiff achieves slightly better transfer attack performance than the original AdvDiff. The reason could be the gradient of the cross-entropy loss is shared among nearly all the deep learning models and is better at attack transferability against these models. But the transferability of AdvDiff can be further improved by selecting different target labels, which is discussed in the Appendix.
@@ 5.5 Ablation Study
Noise sampling guidance steps $N$. Like the iteration times of GAN-based unrestricted adversarial attacks, larger steps $N$ can effectively increase the attack performance against an accurate classifier, as shown in Figure 5. But it can affect the initial noise distribution and hence decreases the generation quality. During experiments, we observe that adversarial guidance is already capable of generating adversarial examples with high ASR. Thus, we can set a small noise sampling guidance step $N$ for better sample quality.
@@
6 Discussions
AdvDiff proves that diffusion models can serve as more effective generation models than GANs for conducting unrestricted adversarial attacks. GAN-based unrestricted adversarial attack methods regularly generate adversarial examples from the target class by optimizing the input noise with the adversarial objective, which is visually perceptible by humans. Our AdvDiff, however, is more feasible to generate unrecognizable adversarial examples from the ground truth label with carefully designed conditional sampling. A potential advantage is that the adversarial examples generated by AdvDiff can be adopted to improve the robustness of deep learning classifiers for achieving both successful attacks and high-quality image synthesis.
@@ 7 Conclusion
In this work, we propose a new method called AdvDiff, which can conduct unrestricted adversarial attacks using any pre-trained conditional diffusion model. We propose two adversarial guidance techniques in AdvDiff that lead diffusion models to obtain high-quality, realistic adversarial examples without disrupting the diffusion process. Experiments show that our AdvDiff vastly outperforms GAN-based attacks in terms of attack success rate and image generation quality, especially in the ImageNet dataset. AdvDiff indicates that diffusion models have demonstrated effectiveness in adversarial attacks, and highlights the need for further research to enhance AI model robustness against unrestricted attacks.

### v4

@@ 4.5 Ablation Study
Noise sampling guidance steps $N$. Like the iteration times of GAN-based unrestricted adversarial attacks, larger steps $N$ can effectively increase the attack performance against an accurate classifier, as shown in Figure 5. However, it can affect the initial noise distribution and hence decreases the generation quality. During experiments, we observe that adversarial guidance is already capable of generating adversarial examples with high ASR. Thus, we can set a small noise sampling guidance step $N$ for better sample quality.
@@ 5 Conclusion
In this work, we propose a new method called AdvDiff, which can conduct unrestricted adversarial attacks using any pre-trained conditional diffusion model. We propose two novel adversarial guidance techniques in AdvDiff that lead diffusion models to obtain high-quality, realistic adversarial examples without disrupting the diffusion process. Experiments show that our AdvDiff vastly outperforms GAN-based and diffusion-based attacks in terms of attack success rate and image generation quality, especially in the ImageNet dataset. AdvDiff indicates that diffusion models have demonstrated effectiveness in adversarial attacks, and highlights the need for further research to enhance AI model robustness against unrestricted attacks.
Acknowledgements
This work was supported in part by HK RGC GRF under Grant PolyU 15201323.